# A General Framework for Joint Multi-State Models

### v3

@@ 5 Dynamic prediction
Let $G=(V,E)$ be a directed acyclic graph, $\mathbb{A}\subset V$ a non-empty subset of states, and $\mathbb{Y}=\mathbb{R}\cup\{+\infty\}$. For any two non-empty sets $\mathbb{A},\,\mathbb{B}\subset V$, we note $\mathbb{A}\rightsquigarrow\mathbb{B}$ if there exists a path from $\mathbb{A}$ to $\mathbb{B}$ in $G$. Let $\tau_{i}=\inf\{n\in\mathbb{N}:S_{in}\in\mathbb{A}\vee\{S_{in}\}\not\rightsquigarrow\mathbb{A}\}$. Then, $\tau_{i}\leq\operatorname{depth}(G)<+\infty$ and take $\chi_{\mathbb{A}}^{*}(\mathcal{T}_{i}^{*})=T_{i\tau_{i}}+\mathds{1}_{\{S_{in}\}\not\rightsquigarrow\mathbb{A}}(+\infty)$ represents the hitting time for the set $\mathbb{A}$.
In Example  1, the stopping time $\tau_{i}$ captures the step at which the process for individual $i$ reaches or exceeds a given time $u$, or enters an absorbing state. The resulting value $\chi^{*}(\mathcal{T}_{i}^{*})=S_{i\tau_{i}}$ therefore represents the state of the individual at that specific time.
@@
This simulation-based dynamic prediction framework generalizes classical approaches from single-event joint models to the multi-state context. It enables individualized forecasting of future state occupancy, event risks, sojourn times, and other clinically relevant outcomes, fully exploiting the subject’s observed biomarker trajectory and event history. In Section 7, we demonstrate its application to forecasting dependency trajectories in the PAQUID cohort.
@@ 6 Simulation Study
The purpose of this simulation study is to evaluate the finite-sample performance and convergence properties of the proposed inference algorithm under controlled conditions where the true parameters are known. We simulate data from a three-state semi-Markov process coupled with a nonlinear longitudinal biomarker trajectory, representing a simplified disease progression model. This setup allows us to assess both statistical accuracy and computational scalability of the stochastic-gradient estimation procedure.
All simulations, estimations, and figures reported in this section were produced using the open-source jmstate Python package, available on PyPI, which we developed to implement the multi-state joint modeling framework and inference algorithms described in Sections 3.2 and 4.
@@ 6.1 Model Specification
The simulated model involves three states—Healthy (0), Sick ($1$), and Terminal (22)—connected by the transitions $0\rightarrow 1$, $0\rightarrow 2$, and $1\rightarrow 2$ (Figure 3). This configuration captures the typical monotone evolution of a disease process, with both direct and indirect paths to the terminal state. Transition intensities follow an exponential baseline hazard under a clock-reset convention.
The longitudinal process follows a continuous piecewise-affine regression:
$h(t,\psi)=\psi_{1}+\psi_{2}t+\mathds{1}_{t>\tau}(\psi_{3}-\psi_{2})(t-\tau),$
with a known inflection time $\tau=6$. The individual-specific parameters $\psi_{i}=(\psi_{i1},\psi_{i2},\psi_{i3})$ arise from a nonlinear mixed-effects model with Gaussian random effects $b_{i}\sim\mathcal{N}(0,Q)$ and independent Gaussian noise $\epsilon_{ij}\sim\mathcal{N}(0,R)$. The link function combines the current biomarker value and its slope,
$g(t,X_{i},\psi_{i})=\big(h(t,\psi_{i}),\frac{\partial h}{\partial t}(t,\psi_{i})\big),$
shared across transitions. Covariates $X_{i}$ are normally distributed and enter linearly through $\beta^{k^{\prime}\mid k}X_{i}$.
Longitudinal data were collected at up to $20$ random time points per subject and censored at $C_{i}\sim\mathcal{U}([10,15])$. Censoring also truncates the biomarker trajectories. The true parameter vector lies in $\mathbb{R}^{16}$. Each simulated dataset comprised $n=1000$ individuals, generating approximately balanced transition counts (Figure 4).
This trajectory mimics a biomarker whose evolution slows after treatment or onset of symptoms, while preserving continuity at the change point.
Figure 3: State transition diagram of the simulated three-state model.
6.2 Estimator Convergence and Accuracy
Parameter estimation was performed using the stochastic-gradient inference procedure of Section 4, with the Adam optimizer (learning rate $0.5$, batch size $K=100$) and an adaptive stopping criterion based on the first and second moments of parameter differences. Convergence was typically achieved in 400400500500 iterations and required around $12$ seconds per run on an AMD Ryzen 5 processor.
Across $n_{\text{runs}}=100$ independent replications, the mean relative bias of all parameters remained below $2\%$, and the root-mean-square error (RMSE) scaled approximately as $n^{-1/2}$. Parameters governing the biomarker trajectory $(\gamma,Q,R)$ were consistently well identified, while association coefficients $(\alpha^{k^{\prime}\mid k},\beta^{k^{\prime}\mid k})$ exhibited slightly higher variability for rare transitions. Representative convergence trajectories are shown in Figure 10, and quantitative results are summarized in Table 5.
Furthermore, longitudinal measurements were taken at $m_{i}\leq 20$ time points, and censoring times were drawn from $C_{i}\sim\mathcal{U}\left([10,15]\right)$. Longitudinal observations were also truncated at the censoring times.
A short summary of the longitudinal process as well as the trajectories is given by the Figure 4 below.
From $\rightarrow$ To # Observations
$0\rightarrow 1$ 613
$0\rightarrow 2$ 375
$1\rightarrow 2$ 592
@@ 6.2 Estimator Convergence and Accuracy
Figure 4: Simulated data: on the left, a sample of longitudinal measurements from $50$ individuals; on the right, observed transitions between states for the complete population of $1000$ individuals.
@@ 6.3 Estimator Convergence
To illustrate the convergence of the estimator, we examine the optimization process for a particular run with $n=1000$ (see Figure 10), as well as RMSE values computed from $n_{\text{runs}}=100$ independent runs on different simulated datasets, all generated from the same underlying distribution. The results of these simulations are shown in Table 5.
To fit the model, we first have to define initial parameters, which can be particularly important for the optimization process if the model is not well-behaved. In the present case, without prior knowledge of the true parameters, we zero the values and use the identity matrix for both covariance matrices.
Formally:
$m_{i}^{(t)}\leftarrow\beta_{i}m_{i}^{(t-1)}+(1-\beta_{i})(\theta^{(t)}-\theta^{(t-1)})^{i},\quad\hat{m}_{i}^{(t)}=\frac{m_{i}^{(t)}}{1-\beta_{i}^{t}},$
and the optimization process is terminated when
$|\hat{m}_{1}^{(t)}|\leq 10^{-6}+10^{-1}\sqrt{\hat{m}_{2}^{(t)}}.$
The Figure 10 shows the evolution of the parameters during the optimization process, and Table 5 summarizes the accuracy of the inferred parameters.
Parameter True value Mean Standard error RMSE
$\gamma_{1}$ 2.5000 2.4960 0.0370 0.0372
$\gamma_{2}$ -1.3000 -1.3039 0.0257 0.0260
$\gamma_{3}$ 0.2000 0.1936 0.0267 0.0275
$\tilde{Q}_{1}$ 0.2554 0.2243 0.0442 0.0540
$\tilde{Q}_{2}$ 0.8047 0.8009 0.0260 0.0263
$\tilde{Q}_{3}$ 0.6020 0.6001 0.0276 0.0277
$\tilde{R} -$0.2653 -0.2684 0.0083 0.0089
$\alpha_{1}^{1\mid 0}$ -0.5000 -0.49997 0.0486 0.0486
$\alpha_{2}^{1\mid 0}$ -3.0000 -2.9962 0.0800 0.0801
$\alpha_{1}^{2\mid 0}$ -1.0000 -0.9990 0.0847 0.0847
$\alpha_{2}^{2\mid 0}$ -5.0000 -4.9897 0.1191 0.1196
$\alpha_{1}^{2\mid 1}$ 0.0000 0.0011 0.0327 0.0327
$\alpha_{2}^{2\mid 1}$ -1.2000 -1.2045 0.0429 0.0432
$\beta^{1\mid 0}$ -1.3000 -1.3015 0.0501 0.0501
$\beta^{2\mid 0}$ -0.9000 -0.8982 0.0670 0.0670
$\beta^{2\mid 1}$ -0.7000 -0.6981 0.0551 0.0551
Figure 5: Comparison of true and estimated parameters (averaged over 100 runs).
The simulation confirms that the proposed inference method accurately recovers all parameters of the multi-state joint model under realistic sample sizes. The optimization demonstrates stable convergence (Figure 10), with minimal sensitivity to initialization. The algorithm’s computational efficiency allows for extensive simulation or cross-validation studies at negligible cost.
@@ 7 Application to the PAQUID Cohort
7.1 Introduction
The data used in this section originates from the PAQUID cohort [Letenneur1994], a large prospective population-based study initiated in southwestern France in 1988, aimed at understanding the determinants and trajectories of aging. A subsample of $n=500$ individuals was followed over a period of up to $20$ years [Proust-Lima2017], with repeated measures of cognitive and physical health, as well as socio-demographic characteristics. In particular, global cognitive functioning was assessed through the Mini-Mental State Examination (MMSE, see Figure 6), while physical dependency was evaluated using the HIER scale, which classifies subjects into four ordered states (see Figure 7): 0 (no dependency), $1$ (mild dependency), $2$ (moderate dependency), and $3$ (severe dependency). In this work, we focus on the association between cognitive decline and the progression of physical dependency by jointly modeling the longitudinal trajectory of MMSE scores and the transitions between the four HIER states using a joint multi-state model. More specifically, the state of an individual $i$ at a given time $t$ is defined as the highest HIER dependency score between trial entry and $t$. This approach allows us to characterize the dynamic interrelationship between cognitive and physical aging processes within a unified statistical framework.
Both time and longitudinal values were normalized in [1,1][-1,1]. This technique ensured stable and fast convergence of the optimization process by keeping all parameters on roughly the same scale.
From $\rightarrow$ To # Observations
$0\rightarrow 1$ 85
$1\rightarrow 2$ 135
$2\rightarrow 3$ 86
$0\rightarrow 2$ 44
$1\rightarrow 3$ 15
$0\rightarrow 3$ 11
Figure 6: On the left, each light gray line represents one of the $500$ observed sequences of (normalized) MMSE scores, and three functional clusters are also represented; on the right, observed transitions between states.
Therefore, the chosen transition graph is as follows, and only allows monotonic transitions from a lower level of physical dependency to a higher one.
@@ 7.1 Introduction
We consider an exponential baseline hazard model for transitions, with a clock reset specification. The parameters of these exponential baseline hazards are jointly optimized with the model parameters.
Moreover, we impose a particular substructure on the model, where $\forall(k,k^{\prime})\in E,\,\beta^{k^{\prime}\mid k}=\beta$.
In the light of Figure 6, the regression and link functions were taken to be scaled and shifted hyperbolic tangent and its derivative respectively such that the normalized MMSE score always has a value of $1$ when $t\to$-$\infty$ and is not increasing
$\begin{cases}h(t,\psi)=\psi_{1}\tanh\left(\frac{\psi_{3}-t}{\psi_{2}}\right)+(1-\psi_{1}),\\ g(t,\psi)=\frac{\partial}{\partial t}h,\\ \psi=\begin{pmatrix}\sigma(\gamma_{1}+b_{1})\\ \exp(\gamma_{2}+b_{2})\\ \gamma_{3}+b_{3}\end{pmatrix},\quad\sigma(z)=\frac{1}{1+e^{-z}}.\end{cases}$
At last, an individual covariates correspond to a pair of binary variables indicating wheter or not the individual has a diploma and is a male $\left(\mathds{1}_{\text{CEP}}(i),\mathds{1}_{\text{male}}(i)\right)\in\{0,1\}^{2}$.

### v4

@@ 5 Dynamic prediction
Let $G=(V,E)$ be a finite directed acyclic graph, $\mathbb{A}\subseteq V$ a non-empty subset of states, and $\mathbb{Y}=\bar{\mathbb{R}}^{+}$. For any two non-empty sets $\mathbb{A},\,\mathbb{B}\subseteq V$, we note $\mathbb{A}\rightsquigarrow\mathbb{B}$ if there exists a path from $\mathbb{A}$ to $\mathbb{B}$ in $G$. Let $\tau_{i}=\inf\{n\in\mathbb{N}:S_{in}\in\mathbb{A}\vee\{S_{in}\}\not\rightsquigarrow\mathbb{A}\}$. Then, $\tau_{i}\leq\operatorname{depth}(G)<+\infty$ and we set $\chi_{\mathbb{A}}^{*}(\mathcal{T}_{i}^{*})=T_{i\tau_{i}}+\mathds{1}_{\{S_{i\tau_{i}}\}\not\rightsquigarrow\mathbb{A}}(+\infty)$ represents the hitting time for the set $\mathbb{A}$.
In Example 1, the stopping time $\tau_{i}$ captures the step at which the process for individual $i$ reaches or exceeds a given time $u$, or enters an absorbing state. The resulting value $\chi^{*}(\mathcal{T}_{i}^{*})=S_{i\tau_{i}}$ therefore represents the state of the individual at that specific time.
@@
This simulation-based dynamic prediction framework generalizes classical approaches from single-event joint models to the multi-state context. It enables individualized forecasting of future state occupancy, event risks, sojourn times, and other clinically relevant outcomes, fully exploiting the subject’s observed biomarker trajectory and event history. In Section 7, we demonstrate its application to forecasting dependency trajectories in the PAQUID cohort (see Subsection 7.3.2).
@@ 6 Simulation Study
The purpose of this simulation study is to evaluate the finite-sample performance and convergence properties of the proposed inference algorithm under controlled conditions where the true parameters are known. We simulate data from a three-state semi-Markov process coupled with a nonlinear longitudinal biomarker trajectory, representing a simplified disease progression model. This setup allows us to assess both statistical estimation error (see Table 5) and computational scalability of the stochastic gradient estimation procedure. We further evaluate the model selection performance of the proposed methodology, as outlined in Subsection 4.2. To ensure statistical reliability, we perform $n_{\mathrm{runs}}=100$ independent simulations on datasets, comprised of $n=500$ individuals, generated from the same underlying distribution.
All simulations, estimations, and figures reported in this section were produced using the open-source jmstate Python package version 0.15.2, available on PyPI, which we developed to implement the multi-state joint modeling framework and inference algorithms described in Sections 3.2 and 4.
@@ 6.1 Model Specification
The simulated model represents a toy example based on a pharmacological scenario where drug absorption can influence seizure activity and patient stability. The model involves two states: Seizure ($1$) and Stable ($2$), which is absorbing. Transitions occur both from $1\rightarrow 1$ and $1\rightarrow 2$ (Figure 3), reflecting how a patient may either experience repeated seizure episodes or stabilize after treatment. If the drug is well absorbed, it helps mitigate seizure activity and promotes stabilization. Conversely, poor absorption increases the likelihood of recurrent seizures, highlighting the importance of pharmacokinetics in disease control. Transition intensities follow an exponential baseline hazard.
Figure 3: State transition diagram of the simulated recurrent two-state model.
The longitudinal biomarker trajectory is modeled using a bi-exponential function:
$h(t,\psi)=\psi_{1}\left(e^{-\psi_{2}t}-e^{-\psi_{3}t}\right),\quad\psi_{i}=\gamma e^{b_{i}},\quad b_{i}\sim\mathcal{N}\left(0,\mathrm{diag}(Q_{1},Q_{2},Q_{3})\right),$
and the observation noise is assumed Gaussian, independent across individuals and time points. This bi-exponential structure is commonly used in pharmacokinetic and longitudinal modeling (e.g., kerioui2022modelling), although alternative parametrizations are sometimes employed. Incidentally, one of the central points in kerioui2022modelling was the deliberate exclusion of recurrent transitions, regarded as difficult to estimate reliably. We show empirically that such transitions can in fact be estimated within our framework.
The link functions are applied consistently across all transitions and defined as
$g^{1\to 1}(t,\psi)=g^{1\to 2}(t,\psi)=\int_{0}^{t}h(s,\psi)\,ds,$
integrating the current biomarker value, thus providing a cumulative measure. Covariates $X_{i}$ are normally distributed and enter linearly via $\beta^{k\to k^{\prime}}X_{i}$.
Longitudinal data were collected at $20$ equally spaced time points between $t=0$ and $t=15$ per subject, with censoring at $C_{i}\sim\mathcal{U}([10,15])$ alongside the individual trajectories.
Since the base hazard rates $\lambda^{1\to 1}$ and $\lambda^{1\to 2}$ are jointly optimized with the other parameters, the final parameter vector lies in $\mathbb{R}^{13}$, and its values used in simulation are shown in Table 5. Each simulated dataset comprised a moderate number of $n=500$ individuals, generating roughly balanced transition counts (Figure 4). Thanks to careful parametrization, almost all parameters are mapped to the whole real axis unconstrained.
A short summary of the longitudinal process as well as the trajectories is given by the Figure 4 below. Note that unlike traditional joint modeling, the notion of censored transitions does not apply here. If a transition does not occur within a trajectory $\mathcal{T}_{i}$, it is simply omitted from the representation.
@@
Figure 4: Simulated data: on the left, a sample of longitudinal measurements from $50$ individuals; on the right, observed transitions histograms for the complete population. $76$ patients experienced repeated seizure.
@@ 6.2 Estimator Convergence
For each run, model parameters were estimated using the stochastic gradient ascent procedure described in Section 4, using a Metropolis-within-Gibbs sampling algorithm. Although the initial parameter values can strongly influence the optimization process, we set $\gamma^{(0)}=(1,1,1)^{T}$ to prevent division by zero, initialize both covariance matrices $Q$ and $R$ as identity matrices, set $\lambda^{1\to 1}$ and $\lambda^{1\to 2}$ to $1$, and initialize all linear coefficients to zero, thereby avoiding any assumptions about prior knowledge of the parameters.
We used the Adam optimizer alongside an adaptive stopping criterion based on a local linear trend diagnostic applied component-wise to the parameters optimization history. Formally, given a fixed window size $T\geq 2$ and any $t\geq T-1$, let
$\{\theta_{j}^{(s)}\}_{s=t-T+1}^{t}$
denote the last $T$ iterates of the $j$-th component. We consider the linear model with intercept
$\theta_{j}^{(s)}=a_{j}+c_{j}s+\varepsilon_{j}^{(s)},\qquad s\in\{t-T+1,\dots,t\},$
and compute the associated coefficient of determination $R_{j,t}^{2}$. Since in such a simple linear regression, the $R^{2}$ coefficient coincides with the squared correlation between the response and the regressor, small values of $R_{j,t}^{2}$ indicate that the recent evolution of $\theta_{j}$ is not better explained by a linear trend than by a constant model, and is therefore consistent with local stationarity.
The optimization procedure is stopped as soon as
$\frac{1}{\dim\theta^{(t)}}\sum_{j=1}^{\dim\theta^{(t)}}R_{j,t}^{2}\leq\delta,$
where $\delta\in[0,1]$ is a prescribed threshold (in practice $\delta=0.1$). With a relatively high value of $T=100$, this criterion that over the last $T$ iterations, no component displays a statistically meaningful linear drift, so the parameter vector is locally indistinguishable from a constant trajectory. In our findings, it proved to be robust in spite of the stochastic nature of the optimization path, and is scale-invariant.
@@
To illustrate the convergence of the estimator, we examine the optimization process for a particular run in Figure 10, as well as RMSE values computed from $n_{\mathrm{runs}}=100$, as shown in Table 5.
Parameter True value Bias Std. deviation RMSE
$\gamma_{1}$ 2.000 -0.018 0.174 0.175
$\gamma_{2}$ 0.200 0.012 0.120 0.121
$\gamma_{3}$ 1.000 -0.009 0.101 0.101
$\frac{1}{2}\log(Q_{1})$ 0.949 -0.017 0.159 0.160
$\frac{1}{2}\log(Q_{2})$ 1.498 -0.035 0.237 0.240
$\frac{1}{2}\log(Q_{3})$ 1.151 -0.040 0.264 0.267
$\frac{1}{2}\log(R)$ 1.498 0.000 0.009 0.009
$\log\lambda^{1\to 1}$ -1.609 -0.009 0.152 0.152
$\log\lambda^{1\to 2}$ -4.605 -0.035 0.176 0.180
$\alpha_{1}^{1\to 1}$ -1.000 0.007 0.148 0.148
$\alpha_{1}^{1\to 2}$ 0.500 -0.010 0.144 0.144
$\beta^{1\to 1}$ -1.000 0.025 0.153 0.155
$\beta^{1\to 2}$ 0.500 0.018 0.143 0.144
Figure 5: Comparison of true and estimated parameters over $100$ runs with $n=500$ simulated individuals (three decimal places). Since the covariance matrix $Q$ is assumed diagonal, the $\log$-Cholesky reduces to the vector of negative $\log$-standard deviations.
The simulation results indicate that the proposed estimation procedure accurately recovers all parameters of the multi-state joint model for moderate sample sizes. Across the $100$ replications, the biases remain remarkably low, and the RMSE values are small compared to the magnitude of the corresponding parameters. Moreover, the observed variability is probably largely attributable to the intrinsic sampling variability of the true maximum likelihood estimator rather than to optimization error.
The population-level parameters $\gamma$ and the measurement noise parameter $R$ display comparatively lower dispersion, indicating stronger sensitivity for these components within the model. The optimization procedure exhibits stable convergence (Figure 10) and further supports that the interpretation that residual fluctuations across replications are primarily driven by the variance of the estimator itself.
6.3 Model Selection
To assess the performance of the information criteria described in Subsection 4.2, we compared the Akaike Information Criterion and the modified $\mathrm{BIC}\mathcal{I}$. Following the same procedure as in the rest of the current section, we performed $n_{\mathrm{runs}}=100$ independent simulations, generating for each run a dataset of size $n=500$ from the above specified model. For each dataset, we then fitted four competing models:
  1. 1.
    Correct model: the model used to generate the data, with correctly specified parameters.
  2. 2.
    Shared-$\alpha$ model: a simplified model in which the link parameters $\alpha^{k\to k^{\prime}}$ are constrained to be equal across all transitions, i.e., $\alpha^{1\to 1}=\alpha^{1\to 2}=\alpha$.
  3. 3.
    Expanded-$\beta$ model: a slightly more complex model in which three additional normally distributed covariates are added to each individual without any association.
  4. 4.
    Link misspecified: a model in which the link functions are incorrectly specified as the bi-exponential function instead of the cumulative exposure.
For each run, the model with the smallest value of AIC or $\mathrm{BIC}_{\mathcal{I}}$ was selected. Table 6 summarizes the results across the $100$ simulations.
Model AIC selections $\mathrm{BIC}_{\mathcal{I}}$ selections
Correct 69 99
Expanded-$\beta$ 31 1
Shared-$\alpha$ 0 0
Misspecified link 0 0
Figure 6: Number of times each model was selected based on AIC and $\mathrm{BIC}_{\mathcal{I}}$ over $100$ simulation runs.
The results indicate that both information criteria successfully identify the correct model in the majority of cases. However, $\mathrm{BIC}_{\mathcal{I}}$ favors the correctly specified model far more frequently than AIC, aligning with its stronger penalty for model complexity, and is able to correctly identify the model $99\%$ of the times.
Indeed, the increase in the AIC penalty for the expanded model is only $6$ relative to the correctly specified model. Yet the negative log-likelihood is of the order of 2 0002\,000. In such a regime, the additional penalty appears negligible compared with the scale of the likelihood term, so the AIC does not provide a sufficient criterion in this situation. Both the shared-$\alpha$ model and the misspecified link models were never selected by the AIC nor the $\mathrm{BIC}_{\mathcal{I}}$.
Overall, these findings demonstrate that our model selection criteria very effectively distinguish between correctly specified and misspecified multi-state joint models, providing reliable guidance for model choice in finite-sample settings.
@@ 7 Application to the PAQUID Cohort
The data used in this section originates from the PAQUID cohort [Letenneur1994], a large prospective population-based study initiated in southwestern France in 1988, aimed at understanding the determinants and trajectories of aging. A subsample of $n=500$ individuals was followed over a period of up to $20$ years [Proust-Lima2017], with repeated measures of cognitive and physical health, as well as socio-demographic characteristics.
In particular, global cognitive functioning was assessed through the Mini-Mental State Examination (MMSE), while physical dependency was evaluated using the HIER scale, which classifies subjects into four ordered states (see Figure 7): 0 (no dependency), $1$ (mild dependency), $2$ (moderate dependency), and $3$ (severe dependency). Individual covariates correspond to a pair of binary variables indicating whether or not the individual has a diploma and is a male $\left(\mathds{1}_{\mathrm{CEP}}(i),\mathds{1}_{\mathrm{male}}(i)\right)\in\{0,1\}^{2}$.
In this work, we focus on the association between cognitive decline and the progression of physical dependency by jointly modeling the longitudinal trajectory of MMSE scores and the transitions between the four HIER states using a joint multi-state model. More specifically, the state of an individual at a given time $t$ is defined as the highest HIER dependency score between trial entry and $t$, and progression is therefore monotone. This approach allows us to characterize the dynamic interrelationship between cognitive and physical aging processes within a unified statistical framework.
7.1 Preprocessing
Longitudinal values were normalized to the interval $[0,1]$ by dividing each score by its maximum value of $30$, and the measurement times were normalized to represent the elapsed time since age $65$ in decades. This technique ensured stable and fast convergence of the optimization process by keeping all parameters on roughly the same scale.
7.2 Model Specification
The transition graph was chosen as follows in Figure 7, and only allows monotonic transitions from a lower level of physical dependency to a higher one, as some transitions occur so infrequently that they cannot be meaningfully analyzed. We considered Weibull baseline hazards for all transitions, consistent with previous observations by Proust-Lima2017, with a clock-reset (sojourn time) specification. The parameters of these baseline hazards were jointly optimized with the model parameters. It should also be noted that not all patients begin in state 0, thus, in this respect, our model inherently accounts for left censoring.
@@ 7.2 Model Specification
The regression and link functions were taken to be shifted and scaled logistic curve such that the normalized MMSE score is always is non-increasing
[…]
For every transition $(k,k^{\prime})\in E$, we define $g^{k\to k^{\prime}}:=g$. The parameter $\Psi_{1}$ governs the amplitude of the logistic function, $\Psi_{2}$ introduces a time shift, and $\Psi_{3}$ controls the slope at the inflection point.
Moreover, we imposed a specific substructure on the model by setting $\forall(k,k^{\prime})\in E,\,\beta^{k\to k^{\prime}}=\beta$, which helped mitigate issues arising from the small number of observations in some groups.
Additionally, since no prior information was available for the covariance matrix of the latent effects $Q\in\mathbb{R}^{3\times 3}$, we used a full $\log$-Cholesky parametrization of the precision matrix $Q^{-1}$, with six parameters denoted $\tilde{Q}_{1}$ through $\tilde{Q}_{6}$. The full parametrization is essential to capture the potential cross-dependencies between the three latent effects.